Foyert Corp. requires a minimum $6,500 cash balance. Loans taken to meet this requirement cost 2% interest per month (paid monthly). Any excess cash is used to repay loans at month-end. The cash balance on October 1 is $6,500 and the company has an outstanding loan of $2,500. Forecasted cash receipts (other than for loans received) and forecasted cash payments (other than for loan or interest payments) follow.

October November December
Cash receipts $ 22,500 $ 16,500 $ 20,500
Cash payments 24,750 15,500 15,500


Prepare a cash budget for October, November, and December. (Negative balances and Loan repayment amounts (if any) should be indicated with minus sign. Round your final answers to the nearest whole dollar.)

FOYERT CORP.
Cash Budget
For October, November, and December
October November December
Beginning cash balance $6,500 $6,500 $6,500
Cash receipts 22,500 16,500 20,500
Total cash available 29,000 23,000 27,000
Cash payments 24,750 15,500 15,500
Interest on bank loan
Preliminary cash balance
Additional loan (loan repayment)
Ending cash balance
Loan balance
Loan balance - Beginning of month $2,500
Additional loan (loan repayment)
Loan balance - End of month
Answer #1
FOYERT CORP.
Cash Budget
For October, November, and December
October November December
Beginning cash balance $                      6,500 $                      6,500 $                   6,500
Add: Total Cash Receipts $                   22,500 $                   16,500 $                 20,500
Total Cash Available $                   29,000 $                   23,000 $                 27,000
Less:Total Cash Payments $                   24,750 $                   15,500 $                 15,500
Less: Interest on bank loans $2,500*2% = $50 $4,800*2% = $96 $3,896*2% = 78
Preliminary cash balance $                      4,200 $                      7,404 $                 11,422
Additional loan (loan repayment) $                      2,300 $                       (904) $                 (3,896)
Ending cash balance $                      6,500 $                      6,500 $                   7,526
Loan Balance
Loan balance - Beginning of the month $                      2,500 $                      4,800 $                   3,896
Additional loan (Loan repayment) $                      2,300 $                       (904) $                 (3,896)
Loan balance - End of month $                      4,800 $                      3,896 $                            -

You can reach me over comment box if you have any doubts. Please rate this answer